OpenStudy (anonymous):

Rain which falls onto a flat rectangular surface of length 6m and width 4m is collected in a cylinder of internal radius 20 cm. What is the depth of water in the cylinder after a storm in which 1 cm of rain fell?

OpenStudy (mayankdevnani):

Volume of water=LWH=600*400*1=240000cm^3

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

so pi*h*r^2=240000 h= 240000/(pi*100)

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

h= 240000/(pi*100) =2400/pi =763.94cm =7.6m

OpenStudy (mayankdevnani):

ok got it

OpenStudy (anonymous):

Is this the final answer?

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

But the answer written in the book should be 191 not 7.6

OpenStudy (mayankdevnani):

oh sry.. conver m into cm

OpenStudy (mayankdevnani):

convert m into cm

OpenStudy (mayankdevnani):

so 600cm and 400 cm

OpenStudy (anonymous):

So it will be 6*4=24m